# Web vs. LLMs: An Empirical Study of Learning Behaviors of CS2 Students

### v1

@@ 5.1.2. Quiz Performance
Activity Treatment Time Spent (%) Minimum Maximum Mean Locating AI 7% 40% 26% NoAI 7% 38% 22% Learning AI 16% 80% 51% (Example) NoAI 3% 73% 40% Learning AI 0% 40% 15% (Theory) NoAI 2% 52% 17% Using IDE AI 0% 37% 8% NoAI 0% 39% 10% Watching Videos NoAI 0% 56% 10%
5.2. RQ2: Learning strategies
5.2.1. Activity Analysis
We qualitatively coded participant activities according to Table 3, and a distribution of the time they spent on various activities is presented in Table 6. In the NoAI treatment, the data for locating information includes the data for locating websites (20.7%) as well as videos (1.4%).
As the table shows, in both conditions, participants showed a strong tendency to learn by example. P21 eloquently described why they used examples to understand concepts: “With examples, I can try to implement them and by creating variations I can understand the concept better. They are easier to comprehend… not abstract”. Others cited that examples were helpful to understand the applications and usefulness of a concept: ”It wasn’t very obvious why currying was useful, but through the examples it is very clear” (P25).
Likewise, in the NoAI treatment, 25% of all 32 participants watched videos, finding that videos were more accessible than websites and made learning easier. For example, P22 said: ”I felt like these websites are directly jumping to the output… I am not familiar with the syntax so it would be good if I can see the flow, and a YouTube video can help with that.”. P28 said that they used YouTube as videos can provide a “demonstration” of the concepts and P20 mentioned that videos felt more like “human interactions” than websites.
Nearly two-thirds of participants (23 out of 32) chose to try examples in the VSCode IDE while learning, across the AI and NoAI conditions (Table 6). However, this amounted to only about 10% of their time on average (Table 5). This can be attributed to the fact that many participants only used the IDE to run code from external sources and verify its expected output, and did not tinker with them. Of all 32 participants, 7 mentioned that they used the IDE only to verify ChatGPT-generated code and 5 mentioned the same for websties.
{mdframed}
[linewidth=0.5pt, linecolor=gray!50, backgroundcolor=gray!10]
Observations: Participants preferred learning by example across both treatments. They did not spend much time coding while learning, often using the IDE only to verify external code. Some participants watched videos while learning as a more dynamic alternative to websites.
5.2.2. Query Analysis
Participants wrote a total of 148 search queries (Google + YouTube) in the NoAI treatment and a total of 237 prompts to ChatGPT. We coded these queries based on the classification scheme in Table 4. Table 5 lists these code occurrences in the web and ChatGPT conditions.
Participants began their learning episodes based on the learning objectives provided to them. Nearly half of all participants (15 out 32) used them directly as the starting point for prompting in the AI condition, in contrast to only 5 out of 32 in web search (NoAI). Most (18 out of 32) participants used keywords based on the learning objectives as the starting point for their web information foraging (code: Statement).
@@ 5.2.2. Query Analysis
Drilling further into the differences in learning behaviors between the AI and NoAI environments, we evaluated the effects of learning environment and phrasing type (Table 4; query phrasing) on the number of queries/prompts issued by participants. A two-way ANOVA of the number of queries (without replication, with repeated measures for each participant) revealed significant effects for learning treatment (F(1, 496) = 16.49, p ¡¡ 0.01), query phrasing, (F(7, 496) = 35.81, p ¡¡ 0.01) and interaction between learning treatment and query phrasing (F(7, 496) = 19.84, p ¡¡ 0.01). Thus, participants preferred some phrasing styles over others in general (Table 4), but also phrased their queries differently in different environments.
The significant interaction effect from the ANOVA test suggests that specific learning environments might promote or discourage certain ways of query phrasing more than others. Table 5 hints at these specific instances of query phrasing differences across the two treatments. For example, participants used more explanation-soliticing (Explanatory) queries in the AI treatment, while in the NoAI treatment they preferred using keyword-based queries (Statement), suggesting a preference for more generic than specific information seeking in the web. This also reflects in the use of the word “example” in only 7% of web search queries as opposed to in 20% of AI prompts. In both conditions, however, participants spent high amounts of time learning by examples (Table  6). Similarly, 57% of LLM prompts were phrased as questions, as compared to 47% in search queries. During a retrospective interview, P31 explained their reason for phrasing queries as questions to ChatGPT but not to Google: “Google doesn’t behave like ChatGPT…. instead of saying ‘what is this’, it’s better to just write the key terms for Google”.
@@ 6. Threats to validity
Internal. We randomized the order of tasks and treatments to minimize learning effects. We had two tasks with different difficulty levels and this stimulated diversity in learning tasks.
Construct. We used quizzes and debugging tasks to evaluate participant learning, which are insufficient. However, we validated our questions and the debugging task in pilot studies to evaluate participants’ knowledge on key concepts in the learning goals.
External. Our results are based on a small sample size relative to the general student population. IIT Kanpur has a female:male ratio of 20%, and despite our best efforts we could not recruit a gender balanced sample. Only  20% of our participants were female. Replication of this study with a gender-balanced sample is needed to generalize our results to other environments. We also need further studies to generalize our findings to other kinds of learning tasks and situations (e.g., concepts in computer architecture or usability).
@@ 7. Discussion and Concluding Remarks
In this paper, we provide a comparative study on how students used ChatGPT and Google to find material and learn new topics in programming in a self-guided manner. As ChatGPT and similar AI tools gain popularity, examining them as learning tools in the context of traditional online learning methods helps us gain insights into learners’ interactions with them.
@@
ChatGPT and other LLMs can improve students’ productivity in performing programming tasks [14], but their effectiveness in helping students learn and understand new concepts in unclear. In our study, while AI helped participants slightly more than web search in successfully debugging programs, a notable number of participants in the AI condition lacked understanding of their proposed fixes (Figure 2). Similarly, participants in the AI treatment performed poorly in the quiz for currying, the more difficult task, as compared to participants in the NoAI treatment, suggesting that the nature of productivity may be at odds with learning. While productivity emphasizes on the speed in completion of tasks, successful learning can be slow and tedious, and ChatGPT might not be the best tool for learning, especially in the absence of sufficient metacognitive abilities [5, 23].
@@
Learning theories have emphasized the importance of examples in learning. Examples can provide learning opportunities for learners to imitate experts, discern analogies and contrast among different problems. In our study, participants across treatments showed a tendency to learn through examples – participants in the AI treatment spent 51% of their time learning by example while those in the NoAI treatment spent 40% of the time learning by example and an additional 10% learning through videos. This emphasizes the importance of incorporating high-quality examples in pedagogy.
@@
Participants in our study tended to write keyword-based queries when using search engines, which led them to websites containing human-curated content about programming concepts. While using LLMs, participants wrote more questions and explicitly asked for information in their prompts. Since LLMs reply directly to these prompts, their responses might only cover the portion of information about the concept that is asked for in the prompt. Our results suggest that students are more likely to ask follow-ups and thus dive deeper into areas they have already explored previously while using LLMs. This might lead learners down a rabbit hole, preventing them from gaining a more complete understanding of the concept, thus explaining participants’ poorer performance in AI condition for the harder task. Future research can explore how LLMs can understand student intent and provide more holistic responses rather than generating direct responses to prompts for exploratory learning. One approach would be to draw inspiration directly from web searches that afford learning from multiple perspectives by utilizing multiple LLM responses. In our study, while using web search, participants often visited more than one website for one search query (211 websites + videos, from 148 search queries).
@@
Modern LLMs generate high-quality textual responses to user prompts. They can provide examples as well as high-quality explanations of these examples[18]. However, participants in the NoAI treatment also expressed interest in non-textual learning resources, notably videos. Some participants chose to watch videos as they are more dynamic than text-based responses, and allow them to better understand the process behind using the concept. Others cited them to be more human. These preferences are consistent with prior work by Jackman et al. [13] on the benefits of YouTube as a tool to demystify complex concepts. Thus, when choosing generative AI tools as learning resources in educational settings, it is imperative to choose ones that allow for heterogenous content types for different learning preferences and styles. For example, [1] can augment its generated textual content with links to video resources. However, there are opportunities for research and development in multimodal language models that can help in creating more effective learning content (e.g., videos, interactive visualizations) and human-like experiences for diverse subjects and learners[17].

### v3

@@ 5.1.1. Activity Analysis
Observations: Participants showed a tendency to seek examples across both treatments. They did not spend much time programming during the learning task, often using the IDE only to verify external code. Some participants preferred to watch videos as a more dynamic alternative to websites.
Code Definition Example
Query Origin
Copied Queries directly pasted from external sources without modification Copying ”What is currying in programming?” from the learning objectives
Edited Queries edited or written completely by participants Modifying the original query to ”What is currying in javascript?”
Follow-Up
Base Queries starting a new line of thought independent of previous queries or learned information What is currying in javascript?
Derived Queries associated with a previous query or learned information that continue exploring an existing line of thought, such as follow-ups, modifications, or clarifications Explain this example in more detail
Query Phrasing
Evidence-Based Queries that ask about the features/description/definition of a concept/idea/object/event What is currying in programming?
Reason You want to find out reasons of/for something. Why are IIFEs used?
Comparison Queries that ask to compare/contrast two or more things, understand their differences/similarities. How are curried functions better than normal functions?
Instruction Queries to understand the procedure/method of doing/achieving something. How do I write an IIFE in javascript?
Debate Queries that debate a hypothetical question (is someone right or wrong, is some event perceived positively or negatively?). Do you really think currying can be useful?
Experience Queries that ask for advice or recommendations on a particular topic. Do you recommend using currying?
Keyword-Based Queries not phrased as a question that do not solicit information Currying in javascript
Explanation Queries not phrased as a question that solicit information Please explain how currying helps with function composition
Table 5. Query Classification Codebook
@@
5.1.2. Prompt and Query Analysis
Participants wrote a total of 237 prompts to ChatGPT and a total of 148 search queries (Google + YouTube) in the NoAI treatment. To understand the nature of these prompts and queries, we categorized them across three dimensions: follow-up, origin, and phrasing (Table 5). As web searches often lead to follow-up queries [37], this analysis helps us examine if and how LLM prompts lead to follow-ups. To understand phrasing, we adapted the ‘Expression’ dimension of the TELer taxonomy of prompts by Karmaker et al.[14], dividing prompts and queries into question-style and instruction-style. The taxonomy for question-style phrasing has been adapted from Bolotova et al.’s [6] taxonomy for questions. For instruction-style queries, we introduce two subcategories - Keyword-Based and Explanation, described in Table 5. Based on coding the prompts and queries of 4 participants, the interrater reliability was 89%, 83% and 81% on Cohen’s Kappa [8] for the dimensions of origin, follow-up, and phrasing respectively. Table 6 lists these code occurrences in the web and ChatGPT conditions.
Participants began their learning episodes using the learning objectives provided to them. Nearly half of all participants (15 out 32) used the text of the learning objectives (described in Section 4.4.1) directly as the starting point for prompting in the AI condition, in contrast to only 5 out of 32 in web search (NoAI). Most (18 out of 32) participants used keywords based on the learning objectives (queries phrased as Keyword-Based) as the starting point for their web information seeking.
@@ 5.1.2. Prompt and Query Analysis
Drilling further into the differences in learning behaviors between the AI and NoAI environments, we evaluated the effects of learning treatment and phrasing type (Table 5; query phrasing) on the number of queries/prompts issued by participants. We modeled this count using a Generalized Linear Model (GLM) with a Poisson distribution and a log link function (dispersion parameter = 1.19). We found a significant interaction effect between learning treatment and query phrasing ($\chi^{2}(4,N=320)=115.11,p=<0.001$) and a significant main effect for learning treatment ($\chi^{2}(1,N=320)=27.08,p=<0.001$) using the chi-square test ($N$ indicates the total number of observations, each one corresponding to one learning treatment and query type for one participant). This main effect indicates that participants used a significantly higher number of prompts in the AI treatment than queries in the NoAI treatment, reiterating the more interactive nature of LLM-based learning.
The significant interaction effect suggests that specific learning environments promote or discourage certain ways of phrasing queries more than others. Table 6 hints at these instances of query phrasing differences across the two treatments. For example, participants used more explanation-soliciting (Explanation) queries in the AI treatment, while in the NoAI treatment they tended to use keyword-based queries (Keyword-Based), suggesting a preference for more generic than specific information seeking in the web. This also reflects in the use of the word “example” in only 7% of web search queries as opposed to in 20% of LLM prompts, though participants spent high amounts of time learning by examples in both treatments (Table  4). Similarly, 57% of LLM prompts were phrased as questions, as compared to 47% in search queries. During a retrospective interview, P31 explained their reason for phrasing queries as questions to ChatGPT but not to Google: “Google doesn’t behave like ChatGPT…. instead of saying ‘what is this’, it’s better to just write the key terms for Google”.
Figure 3. Percentage of prompts and queries over time
Across both treatments, participants showed a tendency to ask more queries during the start of the learning process. Figure 3 shows the distribution of the number of queries and prompts entered by participants over the learning session.
@@
5.2. RQ2: How effective are the information-seeking strategies in the web and AI treatments in terms of conceptual understanding and practical application?
To understand the effectiveness of participants’ information-seeking strategies, we quantitatively analysed participants’ solutions to the quiz and debugging task based on the scoring metric outlined in Section 4.5.
5.2.1. Debugging Task Performance.
[…]
Figure 4. Debugging Task Performance (LoU refers to lack of understanding)
Figure 5. Conceptual understanding Quiz Scores
5.2.2. Quiz Performance
Figure 5 presents the distribution of quiz scores for each concept and learning treatment, scored on a scale of [-8, +8]. We found that the quiz scores for currying ($M=0.66,SD=2.52$) were significantly lower than IIFE ($M=2.69,SD=2.10$), based on the paired t-test ($t(31)=-4.21,p<0.001$). The effect size, measured by Cohen’s d, was $d=-0.86$, indicating a large effect. It also took participants significantly more time to complete the currying quiz compared to IIFE (163.1 vs 97.21 seconds, on average) as per the paired t-test ($t(31)=5.17,p<0.001$). Across all participants and questions in the quiz, participants chose the ”don’t know” option 31 times (15.9% of all options chosen) in the quiz for currying, and 29 times (13.5% of all options chosen) in the quiz for IIFE. This may suggest that currying is a more difficult topic for participants to learn in the study condition than IIFE in general, or that the post-study quiz in currying task was harder than in IIFE. As a result, we compare the NoAI and AI conditions for each task.
AI vs. NoAI in Currying
[…]
AI vs. NoAI in IIFE
[…]
{mdframed}
[linewidth=0.5pt, linecolor=gray!50, backgroundcolor=gray!10]
@@ 6. Threats to validity
Internal. We randomized the order of tasks and treatments to minimize learning effects. We had two tasks with different difficulty levels and this stimulated diversity in information gathering and learning tasks.
Construct. Our assessments involved quizzes and debugging tasks, which are insufficient to evaluate long-duration participant learning and retention. However, we validated our quiz questions and the debugging task in pilot studies to evaluate participants’ knowledge on key concepts in the learning goals.
External. Our results are based on a small sample size relative to the general student population. IIT Kanpur has a female:male ratio of 20%, and despite our best efforts we could not recruit a gender balanced sample. Only  20% of our participants were female. Replication of this study with a gender-balanced sample is needed to generalize our results to other environments. There is also a need for further studies to generalize our findings to other kinds of learning tasks and situations (e.g., concepts in computer architecture or usability).
@@ 7. Discussion and Concluding Remarks
In this paper, we provide a comparative study on how students used AI (ChatGPT) and traditional web resources (Google search and YouTube) to gather information on new topics in programming in a self-guided manner. As ChatGPT and similar AI tools gain popularity, examining them as learning tools in the context of traditional online learning methods helps us understand their efficacy and applicability in different learning scenarios.
@@
ChatGPT and other LLMs can improve students’ productivity in performing programming tasks [15], but their effectiveness in helping students learn and understand new concepts is unclear. In our study, while AI helped participants slightly more than web search in successfully debugging programs, a notable number of participants in the AI condition lacked understanding of their proposed fixes (Figure 5). Similarly, participants in the AI treatment performed poorly in the quiz for currying, the more difficult task, as compared to participants in the NoAI treatment, suggesting that the nature of productivity may be at odds with learning. While productivity emphasizes on the speed in completion of tasks, successful learning can be slow and tedious, and ChatGPT might not be the best tool for learning, especially in the absence of sufficient metacognitive abilities [5, 30].
New Tools, New Challenges
From Table 4, we can observe that the average time participants spent in the AI and NoAI treatment towards programming (8% vs. 10%), locating information (26% vs. 22%), and learning information (66% vs. 68%) are similar. Despite the fact that web search often requires users to navigate through multiple information sources [36] and that ChatGPT can respond directly to user queries to generate explanations and examples, participants spent more time locating information with ChatGPT. This could be due to information-seeking tasks specific to the AI treatment such as thinking of and writing prompts to express intent [27] and scrolling through the chat history to find information from previous responses. Analysing which learning scenarios might be best suited to address the unique challenges of these different information-seeking setups can help define the role of LLMs as learning tools for computing education.
@@
Learning theories have emphasized the importance of examples in learning. Examples can provide learning opportunities for learners to imitate experts, discern analogies and contrast among different problems. In our study, participants across treatments showed a tendency to learn through examples – participants in the AI treatment spent 51% of their time learning by example, while those in the NoAI treatment spent 40% of the time learning by example and an additional 10% learning through videos. This emphasizes the importance of incorporating high-quality examples in pedagogy.
@@
Participants in our study tended to write keyword-based queries when using search engines, which led them to websites containing human-curated content about programming concepts. While using LLMs, participants wrote more questions and explicitly asked for information. Since LLMs reply directly to user prompts, their responses might only cover the portion of information about the concept that is asked for in the prompt. Our results suggest that while using LLMs, students are more likely to ask follow-ups and thus dive deeper into areas they have already explored previously. For the more difficult task, 51% of participants’ queries in the NoAI treatment were keyword-based, as compared to less than 1% for participants in the AI treatment (Table 10). Using more specific prompts might lead learners down a rabbit hole, preventing them from gaining a more complete understanding of the concept, potentially contributing to participants’ poorer quiz performance in the AI treatment for this task. Future research can explore how LLMs can understand student intent and needs to provide more holistic responses rather than generating direct responses to prompts in exploratory learning scenarios. One approach would be to draw inspiration directly from web searches that afford learning from multiple perspectives by utilizing multiple LLM responses.
@@
Modern LLMs generate high-quality textual responses to user prompts. They can provide examples as well as high-quality explanations of these examples [20]. However, participants in the NoAI treatment also expressed interest in non-textual learning resources, notably videos. Some participants chose to watch videos as they are more dynamic than text-based responses, and allow them to better understand the process behind using the concept. Others cited them to be more human. These preferences are consistent with prior work by Moghavvemi et al. [24] on the effectiveness and popularity of YouTube as a tool for information seeking. Thus, when choosing generative AI tools as learning resources in educational settings, it is imperative to choose ones that allow for heterogenous content types for different learning preferences and styles. For example, [1] can augment its generated textual content with links to video resources. However, there are opportunities for research and development in multimodal language models that can help in creating more effective learning content (e.g., videos, interactive visualizations) and human-like experiences for diverse subjects and learners [18].
Appendix
Appendix A Conceptual Understanding Quiz
Tables 7 and 8 display the questions participants answered in the theory assessment for the Currying and IIFE learning task respectively. The correct options are filled in ($\blacksquare$).
Question Options
Passing arguments (for full application of the function) to a curried function and a non-curried function is done with the same syntax. $\square$ Don’t Know $\square$ Yes $\blacksquare$ No
The curried version of a function that takes one argument is the same as the non-curried version. $\square$ Don’t Know $\blacksquare$ Yes $\square$ No
Curried functions always return values (variables, constants, string, etc.) when they are called with any number of arguments. $\square$ Don’t Know $\square$ Yes $\blacksquare$ No
Every nested function is an example of currying. $\square$ Don’t Know $\square$ Yes $\blacksquare$ No
Which of the following are the advantages of Currying a function? (Select one or more)
$\square$ Don’t Know $\blacksquare$ We can create functions with partial functionality of the original curried function. $\square$ The curried version of a function can return the final value with a fewer number of arguments. $\blacksquare$ We can avoid passing the same variable many times to a function by currying. $\square$ We can add extra functionality by currying that is not present in the non-curried function.
Table 7. Currying Conceptual understanding Quiz
Question Options
The variables created within the body of an IIFE can be accessed outside the IIFE. $\square$ Don’t Know $\square$ Yes $\blacksquare$ No
IIFEs can never have a name. $\square$ Don’t Know $\square$ Yes $\blacksquare$ No
IIFEs can be called later in code after they are defined. $\square$ Don’t Know $\square$ Yes $\blacksquare$ No
We can pass arguments to IIFEs. $\square$ Don’t Know $\blacksquare$ Yes $\square$ No
Which of the following are the advantages of IIFEs? (Select one or more)
$\square$ Don’t Know $\blacksquare$ It allows us to reduce the number of global variables $\square$ It is always invoked immediately whenever the compiler sees it. $\blacksquare$ It allows us to isolate a particular part of code functionality. $\square$ It can have extra functionality that cannot be done with traditional functions.
Table 8. IIFE Conceptual understanding Quiz
Appendix B Debugging Assessment
Figures 6 - 9 show the problem statement and relevant snippets of the expected solution for the debugging assessment for the currying and IIFE learning tasks respectively. Participants did not necessarily have to produce code identical to the expected solution, but had to reach the desired output as specified in the problem statement without hardcoding the solution. The results were manually verified by the researcher conducted the study.
// This code uses an IIFE to print the sum of a and b.
// Why isnt it working? Edit the code to fix it, keeping the structure the same.
var a = 20;
var b = 40;
//Start changes below here!
var result;
function adder(){
(function(a, b) {
var result = a + b;
})(a, b);
console.log(result);
Figure 6. IIFE Debugging Task - Problem Statement
//Start changes below here!
var result;
function adder(){
(function(a, b) {
result = a + b;
})(a, b);
adder();
console.log(result);
Figure 7. IIFE Debugging Task - Expected Solution (Snippet)
// This code is supposed to print the values of a, b, and c once, and then print the sum on the next line.
// To clarify, the expected output is:
// a fixed! Value of a:
// 1
// b fixed! Value of b:
// 2
// c fixed! Value of c:
// 3
// 6
// Currently, the code is not working. But you can use the fact that sum is a curried function to fix it!
// Edit the code to make it work as intended without changing the code up to line 30.
//Tip: Take a couple of minutes first to understand the function below
//Do not change
function sum(a) {
console.log(’a fixed! Value of a: ’)
return (b) => {
console.log(’b fixed! Value of b: ’)
return (c) => {
console.log(’c fixed! Value of c: ’)
return a + b + c
var a = 1;
var b = 2;
var c = 3;
//Start changes below here!
suma = sum(a)(b)(c)
console.log(a)
sumb = sum(a)(b)(c)
console.log(b)
sumc = sum(a)(b)(c)
console.log(c)
console.log(sumc)
Figure 8. Currying Debugging Task - Problem Statement
//Start changes below here!
suma = sum(a)
console.log(a)
sumb = suma(b)
console.log(b)
sumc = sumb(c)
console.log(c)
console.log(sumc)
Figure 9. Currying Debugging Task - Expected Solution (Snippet)
Appendix C Currying Task Qualitative Data
Table 9 and Table 10 show the data for participant activities and prompts/queries respectively for the currying learning task.
Activity Treatment Average Time Spent (%)
Locating AI 24%
NoAI 24%
Learning AI 56%
(Example) NoAI 41%
Learning AI 12%
(Theory) NoAI 13%
Using IDE AI 7%
NoAI 6%
Watching Videos NoAI 16%
Table 9. Proportion of time spent by participants in the currying learning task across AI and NoAI treatments
Property Web queries in NoAI Prompts in AI
Total 69 108
Follow-up
Base 72% 56%
Derived 28% 44%
Query Origin
Copied 14% 14%
Edited 86% 86%
Query Phrasing
Evidence-Based 27% 43%
Reason 8% 12%
Instruction 6% 4%
Comparison 1% 3%
Debate 0% 0%
Experience 0% ¡1%
Keyword-Based 51% ¡1%
Explanation 7% 37%
Table 10. Prompt and Query Analysis across participants for the currying learning task